The HEASARC welcomes your participation in a brief survey to capture how users access and utilize HEASARC data, software, and services. The outcome(s) of this survey will be used to guide, prioritize, and plan our activities and development in the coming years. It contains 18 questions, generally takes just a few minutes to complete, and your answers will remain totally anonymous. We thank you in advance for your valuable feedback.
The Neil Gehrels Swift Observatory

Swift Big Events

Timeline of notable spacecraft events that affect data analysis. (Note: this page is a temporary source of information, and will be replaced at a future date with a more complete resource.)

October 31, 2015 BAT triggering disabled

Triggering was disabled as of 21:10 UT on October 31, 2015, (GCN Circ. 18562) and resumed at 17:05 UT on November 6, 2015 (GCN Circ. 18610). Normal BAT operation was achieved at 14:23 UT on November 13, 2015.

December 20, 2012 Malindi ground station out of service

The azimuth drive on the ASI antenna broke on Dec. 20. Service was restored on Feb. 8, 2013, using the U. of Rome Malindi antenna. The volume of Swift telemetry was reduced during the outage including many intervals when the BAT survey data had only 20 energy bins instead of the usual 80.

July 12, 2009 Spacecraft recorder anomaly

An anomaly in the spacecraft data recorder caused the loss of almost all data for observations between about 15:32 UT on July 12, 2009, and about 15:00 UT on July 15, 2009. The anomaly was reported in GCN Circ. 9664, and the resumption of normal operations was reported in GCN Circ. 9669.

August 11, 2007 Gyro Failure

One of the Swift gyros failed at 07:26 UT on August 11, 2007 (GCN Circ. 6760), which caused all three instruments to be powered off and the spacecraft to go into Sun reference pointing. Normal science operations were restored in a series of incremental steps. Distribution of BAT trigger notices resumed Sept. 12th (GCN Circ. 6781), autonomous slewing to new GRBs during normal working hours was enabled on Sept. 24th (GCN Circ. 6825), autonomous slewing was fully restored on Oct. 19th (GCN Circ. 6946),

March 8, 2007 UVOT AT Sequence Changed

The UVOT AT sequence was changed from the values previously set on March 15, 2006. The new sequence is "release13". See the UVOT AT and PT sequences page for further details.

March 7-8, 2007 UVOT PTs were disabled

The observation of UVOT Pre-planned Targets (PTs) was disabled from approximately March 07, 20:00 UT to March 08, 20:00 UT. A list of affected observations will be linked here, when it becomes available.

November 22, 2006 XRT Boresight Changed

A new boresight for the XRT was implemented at 00:10 UT. For details see GCN Circular 5872, an XRT team update on the boresight change.

November 10-22, 2006 Error in Reported UVOT Hardware Window

From 2006-November-10 T22:48:14 to 2006-November-22 T19:57:00 the UVOT was operating with a reduced hardware window, but not reporting this window size correctly in the telemetry. This ultimately affects the fluxes calculated from the UVOT rates. See this note on correcting UVOT frame times for details, including the proper technique for analyzing these data.

March 15, 2006 UVOT AT Sequence Changed

The UVOT AT sequence was changed from the values previously set on Feb. 23, 2006. The new sequence is "release12_2". See the UVOT AT and PT sequences page for further details.

February 23, 2006 UVOT AT Sequence Changed

The UVOT AT sequence was changed from the values previously set on Jan. 10, 2006. The new sequence is "release12_1". See the UVOT AT and PT sequences page for further details.

January 10, 2006 UVOT AT Sequence Changed

The UVOT AT sequence was changed from the values previously set on Oct. 3, 2005. The new sequence is "release12". See the UVOT AT and PT sequences page for further details.

October 3, 2005 UVOT AT Sequence Changed

At 15:59 UTC the UVOT AT sequence was changed from the launch values to "release11_2". For the original and new sequences, see the UVOT AT and PT sequences page.

September 14, 2005 UVOT Shift-and-Add Change

UVOT changed the "shift-and-add" algorithm used on-board the satellite for aspect correction; the change corrected some losses ascribed to the earlier algorithm.

June 9, 2005 UVOT Flight Software Change

On June 9, 2005, the UVOT team installed a patch to the flight software in order to (1) increase the size of the GeNIe image and (2) allow UVOT to correctly use the XRT position for afterglows.

May 10 - June 21, 2005 BAT TDRSS lightcurves unreliable and/or suspended

A bug was found in the BAT flight software that results in bad times for all BAT lightcurves generated for TDRSS, dating from 17:29 on May 10, 2005. (See GCN circ #3429.) BAT lightcurves calculated on the ground (ie., not in the "tdrss" directory or from TDRSS), are correct. The problem was fixed in a flight software upload at 13:38 on June 7, 2005. Distribution of correct BAT TDRSS lightcurves resumed June 22, 2005. Any BAT TDRSS lightcurves from the period of May 10 - June 7 should be disregarded.

April 7, 2005 UVOT Settling Exposure Change

At 18:55:43 on Apr. 7, 2005, new flight software was enabled that changed the UVOT settling exposure from UVW2 to V.

March 31, 2005 Star-Trackers Lost Lock - Bad Attitude period

Near approx. 09:30-09:40 UTC on March 31, star cameras lost lock, leading to apparent bursts caused by Vela X-1 being detected at a wrong location. Valid attitude was regained at approx. 21:30 UTC. All data between these times have invalid attitude.

March 5 - 29, 2005 UVOT Drift Corrections Enabled On-Board
Date Action
Mar. 5, 2005 UVOT preliminary drift correction enabled.
(Effects UVOT image mode data; preliminary drift correction with this algorithm is incomplete.)
Mar. 7, 2005 UVOT drift correction roll offset parameter updated.
(Improves UVOT image mode data, but drift correction is still problematic during this period.)
Mar. 24, 2005 UVOT drift correction disabled.
Mar. 29, 2005 New UVOT drift correction parameters uploaded to fix drift correction. On-board drift correction re-enabled.
March 9 - 13, 2005 BAT High Voltage Changes and Reboot
Date and Time Action
Mar. 9, 2005 22:15 UTC BAT High Voltage (HV) lowered from nominal 200v to 100v, in 50v steps.
Mar. 10, 2005 15:52-17:58 UTC BAT rebooted to help clear memory
Mar. 11, 2005 18:11 UTC BAT High Voltage (HV) raised from 100v to 200v, in 50v steps, 1 minute apart.
Mar. 11, 2005 19:55 UTC BAT triggers re-enabled
Mar. 13, 2005 01:99 UTC BAT automatic slewing resumed (AT_Slewing Enabled)
February 11, 2005 Period with No XRT Science Data

From 14:49:02 UTC on Feb. 11, 2005 until the end of the day, there is no XRT science data.

January 31, 2005 Stellar Aberration Correction Changed On-board Satellite

At 21:57 UTC, Jan. 31, 2005 the star tracker stellar abberation correction was changed; sources on the focal plane now will have fixed positions in raw coordinates.

December 13, 2004 Star Tracker Boresight Adjustment

At 01:55 UTC on day-of-year 348 (December 13, 2004) a command was uploaded to align the star-tracker boresight with the NFI instrument axes.